OpenStudy (anonymous):

( 12^(x+1))/8^(x+2) = p(n)^ x find that p and n

OpenStudy (anonymous):

someone pls help me ...

OpenStudy (badhi):

\[\begin{align*} \frac{16^{(x+1)}}{8^{(x+2)}} &= \frac{16 \times 16^x}{8^2 \times 8^x} \\ &= \left(\frac{16}{8^2}\right)\left(\frac{16}{8} \right)^x\\ \end{align*}\] Is this helpfull?

OpenStudy (anonymous):

thank

OpenStudy (badhi):

you are welcome :)